an unknown gas occupies 10dm3 of space at 4 atm and 20°c. what is the amount in moles of the gas present if the gas is a real gas? take R = 0.082
1
Expert's answer
2020-03-05T08:18:20-0500

PV=nRT

P = 4 atm = 405 kPa

T = 293,15 oK

n= PV/RT = (405×10)/(8.2×293.15) = 1.68 mol

Answer: 1.68 mol.